OpenStudy (anonymous):

y=x^2+7x+12 I need to find 2 answers

OpenStudy (angela210793):

D=b^2-4ac=49-4*1*12=1 X1=(-b-sqrD)/2a=-7-7/2=-7 X2=(-b+sqrD)/2a=-7+7/2=0/2=0